OpenStudy (anonymous):

Hello. Given: Simplify the expression. \[\LARGE {(a+b)^2-4ab \over a^2+b^2 }\] Attempted: \[\LARGE \begin{array}{l}\frac{{{{\left( {a + b} \right)}^2} - 4ab}}{{{a^2} + {b^2}}} =\\ \\\\ = \frac{{{a^2} + 2ab + {b^2} - 4ab}}{{{a^2} + {b^2}}} = \\\\ = \frac{{{{(a - b)}^2}}}{{{a^2} + {b^2}}}\end{array}\] I also tried some different ways but my answer doesn't math with multiple choices. What's wrong?

OpenStudy (anonymous):

Here's a photo.

OpenStudy (anonymous1):

Your answer is correct. But there doesn't seem to be a match in the multiple choices.

OpenStudy (anonymous):

that's what I thought too. :( ...

OpenStudy (anonymous):

there is a answer. wait let me try.

OpenStudy (anonymous):

ok.... ;)

OpenStudy (anonymous1):

There is no answer in the multiple choices.

OpenStudy (anonymous):

... Even If we split it up, it doesn't match anyway ... \[\LARGE \begin{array}{l}\frac{{{{\left( {a + b} \right)}^2} - 4ab}}{{{a^2} + {b^2}}} = \\\\ = \frac{{{a^2} + 2ab + {b^2} - 4ab}}{{{a^2} + {b^2}}} = \\\\ = \frac{{{a^2} + {b^2} - 2ab}}{{{a^2} + {b^2}}} = 1 - \frac{{2ab}}{{{a^2} + {b^2}}}\end{array}\]

OpenStudy (anonymous):

@anonymous1 you're right... I just wanted to be sure, ;) Thanks everyone for helping. And I have to confess that my book sucks ! LOL ;)

OpenStudy (anonymous1):

OK. The language of this book is Albanese?

OpenStudy (anonymous):

Yes. :) , Google translate? :P

OpenStudy (anonymous):

Albanian.

OpenStudy (anonymous1):

I just recognized it as Albanian (I don't speak it though).

OpenStudy (anonymous):

then how did you know it ? :O ... Sometimes I think I'm the only one in the planet who knows this language LOL.

OpenStudy (anonymous1):

I have already seen some words in Albanian, but I think I recognized it mostly because of the "ë" (e with two dots above).

OpenStudy (anonymous):

haha.. True :$ . it spells like a bird this "a" it sound the same like "ë" :P ... anyway. ;) Thanks for helping .

OpenStudy (anonymous):

i solved that kind of problem before, although i forgot it now. sorry dude.

OpenStudy (anonymous1):

You're welcome.

OpenStudy (anonymous):

@kumarsajan It's ok. ;) @anonymous1 :)